OpenStudy (anonymous):

Solve by factoring. n^2+2n-24=0

OpenStudy (matt101):

Find the factors that multiply to -24 and add to 2, in this case 6 and -4. Therefore: n^2 + 2n - 24 = 0 (n + 6)(n - 4) = 0 n = -6, 4